Consider the following four-alternative decision problem, where x and y are real numbers: Alternative WTP Explicit Cost Choice 1 Choice 2r- 1 Choice 3r- 2 Choice 4- 3 y+1 Select the correct answer among the choices below. Choice 1 is the best choice Choice 2 is the best choice Choice 3 is the best choice Choice 4 is the best choice The decision maker is indifferent between the four choices
Answer #1

Answer Option (5).

The choice for which (WTP - Explicit cost) is maximized is the best choice.

Choice 1: [x - (y + 3)] = x - y - 3

Choice 2: [(x - 1) - (y + 2)] = (x - 1 - y - 2) = x - y - 3

Choice 3: [(x - 2) - (y + 1)] = (x - 2 - y - 1) = x - y - 3

Choice 4: [(x - 3) - y] = x - y - 3

Since (WTP - Explicit cost) is equal for all options, decision maker is indifferent between all choices.